A biquadratic equation must have at least one real root is it true or false ? Give reason?
1
Expert's answer
2020-02-28T13:10:59-0500

A biquadratic equation must have at least one real root is it true or false ?


A biquadratic equation is a polynomial equation with degree 4 without having degree 1 and 3 terms. Since it is saying it "must have" atleast one real root, it makes the statement false.


To prove this I will provide a counter example.

Consider a biquadratic equation

"x^4+7x^2+12=0"

For finding roots , let "u=x^2"

Given equation will rewritten as

"u^2+7u+12=0"

"u^2+3u+4u+12=0"

"u(u+3)+4(u+3)=0"

"(u+3)(u+4)=0"

i.e. "u= -3" or "u = -4"

since "u=x^2"

"x^2=-3" or "x^2=-4"

i.e. "x=\pm \sqrt{-3}" or "x=\pm \sqrt{-4}"

i.e "x=\pm \sqrt{3}i" or "x=\pm 2i"

Which means a biquadratic equation can have all imaginary roots.